What is the proportion to this word problem?
A motorcycle can travel 705 miles on 19 gallons of gasoline. How many gallons of gas are needed to travel 1253 miles?

Answer by lenny460(1073)

What is the proportion to this word problem?
A motorcycle can travel 705 miles on 19 gallons of gasoline. How many gallons of gas are needed to travel 1253 miles?

19/705 = x/1253
Cross multiply:
19 * 1253 = 23807
705 * x = 705x
Therefore:
705x = 23807
Divide each side by 705.
705x/705 = 23807/705
x = 37.1 gallons
He needs 37.1 gallons to travel 1253 miles.


Lennox Obuong
Algebra Tutor